### v1

# Molecular Communication Channel as a Physical Reservoir Computer

Abstract
Molecular Communication (MC) channels inherently possess significant memory and nonlinear dynamics due to diffusion and receptor kinetics, often posing challenges for reliable data transmission. This work reconceptualizes these intrinsic properties as computational resources by framing a canonical point-to-point MC channel, consisting of ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a physical reservoir computer (PRC). We utilize the time-varying fraction of bound receptors as the reservoir’s internal state, employing time-multiplexing to generate high-dimensional virtual nodes without explicit recurrence. Only a linear readout layer is trained via ridge regression. Through deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s capability for complex temporal processing by successfully performing next-step prediction on standard chaotic time-series benchmarks (Mackey-Glass and NARMA10). Performance, quantified by Normalized Root Mean Square Error (NRMSE), exhibits a non-monotonic dependence on key system parameters (receptor kinetic rates, diffusion coefficient, transmitter-receiver distance), revealing optimal operational regimes. These findings validate the potential of using MC channel as effective and low-complexity computational substrate.
Index Terms:
Reservoir Computing, Molecular Communications, Ligand-Receptor Interactions
I Introduction
Reservoir Computing (RC), introduced independently as the echo‑state network (ESN) by Jaeger [1] and the liquid‑state machine (LSM) by Maass et al. [2], is a class of recurrent neural network models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression. This separation of representation and learning obviates back‑propagation through time, providing fast, data‑efficient training while retaining the rich temporal processing power of recurrent dynamics. As a result, RC has proven highly effective for time-series prediction tasks [3].
Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implements large virtual networks through time-multiplexing of a single nonlinear node [8, 9]. Such implementations greatly reduce hardware complexity while achieving performance on par with digital RC, and have catalyzed interest in neuromorphic edge computing [10] and embodied intelligence [11].
Recent work has begun to evaluate biological systems as reservoirs, capitalizing on their inherently rich and adaptive dynamics. Pioneering experiments with neuron cultures [12], brain organoids [13], bacteria [14], and plants [15] demonstrate that living systems can be co-opted or integrated into artificial computing frameworks to execute externally defined computational tasks. Likewise, chemical reaction networks, which underlie many fundamental biochemical processes, have been investigated for RC, demonstrating the potential to exploit complex molecular interactions for sophisticated information processing beyond their native biological context [16].
Central to the exceptional information processing capabilities of biological systems is biochemical signaling, which enables coordination and adaptation across multiple scales, critical for their survival [17]. Building on these naturally evolved mechanisms, researchers have long studied engineered molecular communication (MC) systems as alternative to conventional electromagnetic (EM) communications to enable networks of micro/nanoscale devices via the exchange of precisely controlled chemical signals [18]. This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19]. Extensive theoretical research and recent experimental work in this field have particularly focused on overcoming the unique challenges of MC channels that are different than EM communications, including significant channel memory effects (inter-symbol interference) and nonlinear receiver responses (involving receptor saturation) in addition to biochemical noise [20, 21]. These complexities turned out to be particularly demanding for small, resource-limited devices, yet they also hint at untapped computational potential that biological systems may be inherently exploiting.
In this paper, we reconceptualize MC channel as a physical reservoir computer, revealing its potential for information processing and implementing machine learning tasks. Specifically, we study a simple MC channel model, which is adequately representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors. By adopting a time-multiplexed RC framework, we leverage the inherent channel memory arising from diffusion and the nonlinear ligand-receptor interactions at the receiver, both of which are critical for computational performance. Through numerical analysis and particle-based spatial stochastic simulations, we demonstrate that MC channels can successfully perform chaotic time-series prediction tasks with minimal training, as benchmarked by standard RC evaluations on NARMA10 and Mackey-Glass time-series. We systematically investigate how various MC system parameters, including the number of transmitted molecules, transmitter-receiver distance, diffusion coefficient, and ligand-receptor kinetic rates, affect computing performance.
Our results provide evidence that even simple MC channels can be harnessed for non-trivial time-series tasks by virtue of their intrinsic diffusion-driven memory and receptor-mediated nonlinearities. Beyond contributing to the fundamental understanding of information-processing capabilities of biological systems at the elemental level of their communication links, this computing perspective on MC channels can open new avenues for bio-inspired and bio-integrated computing architectures. Potential applications range from early disease diagnosis analyzing time-varying biomarker profiles to channel equalization in artificial MC networks and IoBNT [22]. By demonstrating how the channel’s inherent dynamics can be systematically exploited for information processing, this study lays the groundwork for MC systems to serve not only as physical channels for communication but also as powerful computational substrates in their own right.
The paper is organized as follows. Section II reviews molecular communication and physical reservoir computing. Section III formalizes the molecular‑communication channel as a physical reservoir computer and describes the modeling and training methodology. Section IV presents deterministic and stochastic simulation results with parameter studies. Section V concludes and outlines future directions.
II Background
II-A Molecular Communications
Bio-inspired MC emerged as a promising approach to realize networks of micro/nanoscale devices, such as biosensors and microrobots, and to seamlessly interface these engineered networks with biological systems. Such integration underpins the vision of the IoBNT, where MC mediates collaborative tasks such as targeted drug delivery and continuous health monitoring [23]. By exploiting mechanisms similar to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods become impractical [24].
From early theoretical studies on capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [21, 25, 26, 27]. Researchers have identified unique challenges that distinguish MC from EM communications, including severe inter-symbol interference (ISI) due to slow diffusion, and nonlinearity plus memory effects from stochastic receptor binding and other reaction kinetics. Practical MC transceivers at the microscale also face tight constraints in energy, size, and computational resources, motivating new strategies in encoding, detection, and synchronization [20]. To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [28].
In parallel, an expanding line of research investigates MC for in situ artificial intelligence (AI) with living cells and engineered biomolecular processes, commonly referred to as wetware AI. Two complementary strategies have emerged: top-down approaches that leverage existing cellular functionalities, and bottom-up approaches that build computing frameworks from fundamental chemical reactions. In the former category, wet-neuromorphic computing was introduced by repurposing gene regulatory networks (GRNs) for pattern recognition [29], while Wet TinyML concept applied GRN-based methods for energy-efficient neural-like computations within living cells [30]. Another study integrated biologically plausible spiking neurons into reservoir computing, using MC-based connectivity for distributed processing among cells [31]. A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GNRs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [32].
By contrast, bottom-up efforts explicitly engineer MC as the foundation for novel neural network architectures. Nano-scale computing units and Molecular Nano Neural Networks (M3N) have been proposed, wherein diffusion-based molecular signaling implements neural-layer functions for in-body intelligence [33]. Another line of work introduced a broader vision of chemical AI, illustrating how reaction networks, synthetic biology, and nanofluidic systems, underpinned by molecular signaling, could equip synthetic cells with self-contained, biocompatible computation [34].
Altogether, these works highlight the synergy between biological information processing, MC, and ML in paving the way for low-power and adaptive intelligence at micro- and nanoscales and within biological contexts. Here, we introduce a new perspective by reframing an elemental point-to-point channel, comprising diffusion and receptor-ligand interactions, as a physical reservoir computer. Unlike previous studies relying on genetic or multi-cellular networks and specialized biochemical circuits, our focus is on systematically exploiting the channel’s intrinsic memory and nonlinearity to perform non-trivial learning tasks with minimal training overhead. Our findings reveal new pathways for leveraging MC channels as both a communication medium and a computational resource, thereby extending earlier demonstrations of wetware AI.
Figure 1: (a) RC framework for a multiclass classification task, (b) An example of physical RC for binary classification: the wave-interference patterns produce a nonlinear mapping of the data, which is then trained for classification [35], (c) Delay-based RC framework with a singular nonlinear node for multiclass classification task, (d) An end-to-end MC system as a reservoir computer with virtual nodes for time-series prediction task.
II-B Physical Reservoir Computing
In a standard RC scheme, the input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
$\mathbf{x}(t+1)=f\!\bigl{(}\mathbf{A}\,\mathbf{x}(t)+\mathbf{W}_{\mathrm{in}}\mathbf{u}(t)\bigr{)},$ (1)
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinearity (Fig. 1(a)) [36]. The readout is obtained by a linear map
$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t),$ (2)
with $\mathbf{W}_{\mathrm{out}}$ trained, often by ridge regression, to minimize the discrepancy between $\mathbf{y}(t)$ and a target signal. Because the internal weights $\mathbf{A}$ remain fixed, RC sidesteps the heavy cost of back‑propagating gradients through deep recurrent structures, a feature that makes it well‑suited to hardware realizations.
PRC replaces the software‑defined reservoir with the intrinsic dynamics of a physical substrate. Phenomena such as optical feedback, spin‑wave interactions, analogue electronic circuits, and even the interference patterns of water waves in a bucket implement the operations $\mathbf{A}\mathbf{x}(t)$ and $f(\cdot)$ in real time, while external sensors capture the instantaneous state $\mathbf{x}(t)$ (Fig. 1(b)) [9, 35]. Leveraging the medium’s native parallelism, bandwidth, and energy efficiency, PRC offers a neuromorphic alternative to digital simulations.
Time‑multiplexed reservoirs. High‑dimensional states can be generated with minimal hardware by time‑multiplexing a single nonlinear element (Fig. 1(c)) [8]. During each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants. The resulting vector
$\mathbf{x}(t)=\bigl{[}x_{1}(t),x_{2}(t),\dots,x_{M}(t)\bigr{]}^{\!\top}$
acts as an $M$‑node virtual reservoir whose evolution from one period to the next is governed by the masked input and the device’s nonlinear response [37].
A widely adopted form of time‑multiplexing employs an external feedback loop of delay $\tau$ around a single nonlinear node (e.g., a semiconductor laser or an analogue circuit). Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ yields $M$ virtual nodes [9, 38]. Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal supplies fading memory that underpins tasks such as speech recognition and chaotic time‑series forecasting. Additional delay lines or delayed taps at the input or readout can further enrich the dynamics or extend the memory horizon without increasing the number of physical components [39].
III Molecular Communication Channel as a Physical Reservoir
We consider a single-link MC system consisting of a point transmitter and a spherical receiver, as illustrated in Fig. 1(d). The transmitter release a prescribed number of ligands (molecules) instantaneously at prearranged times into a fluidic medium, where they propagate by free diffusion. A fraction of these ligands eventually arrives near the receiver and can bind reversibly to receptors on its surface under given reaction kinetics. Below, we describe how this widely studied MC system model can be reframed as a physical reservoir computer by regarding the time-varying fraction of bound receptors as the high-dimensional dynamical state.
Let the transmitter release $I(n)$ molecules at each symbol time $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval. For simplicity, we assume the receiver is at distance $x_{R}\gg r_{\mathrm{rx}}$ from the transmitter at the origin, such that the receiver’s finite radius $r_{\mathrm{rx}}$ can be neglected in the channel impulse response. Under three-dimensional (3D) free diffusion with diffusion coefficient $D$, the ligand concentration $c(\mathbf{x},t)$ obeys Fick’s second law:
$\frac{\partial c(\mathbf{x},t)}{\partial t}\;=\;D\,\nabla^{2}c(\mathbf{x},t),$ (3)
with the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$. The impulse response at distance $x_{0}$ for a single pulse released at $t=0$ is the 3D diffusion kernel
$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl{(}-\frac{x_{0}^{2}}{4Dt}\Bigr{)}\quad(t>0).$ (4)
Hence, when the transmitter emits $I(n)$ molecules at times $nT$, the local ligand concentration at the receiver (treated as a point in the far-field approximation) is given by
$c_{T}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl{(}t-nT\bigr{)}\,u\bigl{(}t-nT\bigr{)},$ (5)
where $u(\cdot)$ is the unit step function enforcing causality. Since $h(t)$ has a infinitely long tail (though it decays rapidly), successive transmissions introduce ISI. For practical numerical simulations, this tail is often truncated to a finite channel memory length [40].
Figure 2: (a) An instance from particle-based stochastic simulations of the MC channel with Smoldyn. Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey-Glass series input, (c) NARMA10 Series input, (d) Mackey-Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
The spherical receiver is equipped with $N_{R}$ identical, monovalent ligand receptors. Each receptor can be either unbound or bound to one ligand, with reversible reaction:
$L+R\overset{k_{\text{on}}}{\underset{k_{\text{off}}}{\longleftrightarrow}}LR,$
where $L$ is the free ligand, $R$ is an unbound receptor, and $LR$ is a ligand-receptor complex, i.e., bound receptor. At the mean-field level, and assuming the local free ligand concentration is not significantly depleted by bound receptors, the time evolution of the mean fraction of bound receptors, $b(t)=n_{b}(t)/N_{R}$, can be approximated by
$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{T}(t)\,\bigl{[}1-b(t)\bigr{]}\;-\;k_{\mathrm{off}}\,b(t),$ (6)
where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the forward (binding) and reverse (unbinding) rate constants, respectively, and $c_{T}(t)$ is the local ligand concentration at the receiver.
Although the exact time-varying variance of $b(t)$ under such nonstationary input is typically not available in closed form (due to correlated ligand arrivals and receptor occupancy fluctuations), we use particle-based spatial stochastic simulations (with Smoldyn (Fig. 2) [41]) in Section IV to account for this stochasticity. For the present analytical treatment, (6) provides a deterministic approximation to the mean behavior of the ligand-receptor interactions.
[…]
In RC, one aims to transform an input time series $\{u(n)\}$ into a high-dimensional dynamical state $\mathbf{x}_{n}$ through a fixed nonlinear system (the reservoir), and then train only a linear readout for the target computational task. Here, we map the input $u(n)$ to the number of released molecules $I(n)$ at each symbol time $nT$:
$I(n)\;=\;\alpha\,u(n),$ (7)
where $\alpha$ is a scaling factor. The channel then imposes a complex, memory-rich temporal response on these inputs (due to both diffusion and receptor binding).
To extract a sufficiently high-dimensional state from the scalar output $b(t)$, we perform time-multiplexing. That is, within each interval $[nT,\,(n+1)T]$, we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$:
$\mathbf{x}_{n}\;=\;\bigl{[}b\bigl{(}nT+\tau_{1}\bigr{)},\;b\bigl{(}nT+\tau_{2}\bigr{)},\;\dots,\;b\bigl{(}nT+\tau_{M}\bigr{)}\bigr{]}^{\top}.$ (8)
This sampling effectively creates $M$ virtual nodes per symbol. Unlike a typical optoelectronic delay-based reservoir, however, we do not require a separate external feedback loop; rather, the channel’s intrinsic ISI provides the dynamical memory. The “masking” function typically used in conventional time-multiplexed RC is replaced here by the MC channel impulse response, which shapes the input in time to eventually generate diverse node responses upon sampling.
We train a linear readout to map from the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a time-series prediction). Let $\mathbf{W}_{\mathrm{out}}$ be the $M$-dimensional readout weight vector, then
$\hat{y}_{n}\;=\;\mathbf{W}_{\mathrm{out}}^{\!\top}\,\mathbf{x}_{n}.$ (9)
With $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, we solve:
$\min_{\mathbf{W}_{\mathrm{out}}}\Bigl{\|}\mathbf{y}\;-\;\mathbf{X}\,\mathbf{W}_{\mathrm{out}}\Bigr{\|}^{2}\;+\;\lambda\,\|\mathbf{W}_{\mathrm{out}}\|^{2},$ (10)
where $\mathbf{X}$ is the $N_{\mathrm{train}}\!\times\!M$ matrix whose $n$-th row is $\mathbf{x}_{n}^{\top}$, $\mathbf{y}$ is the vector of target outputs $y_{n}$, and $\lambda$ is a regularization parameter (ridge regression). The closed-form solution is then given by the Moore–Penrose pseudoinverse:
$\mathbf{W}_{\mathrm{out}}\;=\;\bigl{(}\mathbf{X}^{\top}\,\mathbf{X}\;+\;\lambda\,\mathbf{I}\bigr{)}^{\!-1}\,\mathbf{X}^{\top}\,\mathbf{y}.$ (11)
Hence, only $\mathbf{W}_{\mathrm{out}}$ is optimized, and the underlying channel and binding dynamics remain unmodified.
IV Results
In this section, we present the results of the performance evaluation for the MC reservoir computer using two complementary approaches. First, deterministic analyses were carried out in MATLAB using the mean‑field kinetic model (Eq. (6)) with the default simulation settings listed in Table I. Second, to capture the stochasticity of the MC channel, we performed particle‑based spatial stochastic simulations in Smoldyn [41]. By comparing the outcomes of the deterministic and stochastic simulations, we both confirm the validity of the approximate deterministic model and quantify the effects of intrinsic noise on RC performance.
TABLE I: Default Values of Simulation and RC Parameters
Parameter Value
Receptor binding constant $k_{\text{on}}$ $6.022\times 10^{8}\,\mathrm{M^{-1}s^{-1}}$
Receptor unbinding constant $k_{\text{off}}$ $1\,\mathrm{s^{-1}}$
Maximum # released molecules $N_{\max}$ $3000$
Minimum # released molecules $N_{\min}$ $100$
Symbol duration $T$ $1\,\mathrm{s}$
Mackey–Glass prediction length $\Delta$ $6$ steps
Diffusion coefficient $D$ $1\times 10^{-11}\,\mathrm{m^{2}s^{-1}}$
Tx–Rx distance $x_{0}$ $10\,\upmu\mathrm{m}$
Reservoir nodes $M$ $100$
Numerical (deterministic) soln. timestep $\Delta t_{d}$ $1\,\mathrm{ms}$
Stochastic simulation timestep $\Delta t_{s}$ $10\,\mathrm{ms}$
Stochastic simulation receiver radius $R$ $3\,\upmu\mathrm{m}$
Training sequence length $2000$
Test sequence length $2000$
IV-A Chaotic Time-Series Prediction
We benchmark our MC reservoir computer on two widely used chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [42]. These tasks have become standards for assessing both algorithmic and physical RC architectures because they combine complex temporal dependencies, strong nonlinearities, and a requirement for long‑range memory. The goal is to learn a mapping from the present (and past) states of the time series to future values, using only a linear readout layer trained on the reservoir states.
Performance Metric: We quantify the predictive accuracy of the reservoir using the normalized root-mean-square error (NRMSE),
$\text{NRMSE}\;=\;\sqrt{\frac{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\hat{y}_{k}\,]^{2}}{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\bar{y}\,]^{2}}},$
where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples. An NRMSE closer to 00 indicates a more accurate prediction.
Figure 4: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey-Glass series, and (d-f) for NARMA10 series prediction tasks.
IV-A1 Mackey-Glass Series
The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
$\frac{dx(t)}{dt}=\beta\,\frac{x(t-\tau)}{1+x(t-\tau)^{n}}-\gamma\,x(t),$ (12)
where $\beta$, $\gamma$, $n$, and $\tau$ are positive constants governing the nonlinear feedback loop. For the canonical parameter set $\beta=0.2$, $\gamma=0.1$, $n=10$, and $\tau=17$, the system exhibits chaotic behavior with a fractal attractor.
We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation using a suitable step size (e.g., $0.1$), then sampling at discrete time points $k$. The reservoir input at time step $k$ is set to $u(k)=x_{k}$, and the task is to predict $x_{k+\Delta}$, i.e., the value of the series $\Delta$ steps into the future.
IV-A2 NARMA10 Series
A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by
$y(k+1)$ $=0.3\,y(k)\;+\;$ (13)
$0.05\,y(k)\sum_{i=0}^{9}y(k-i)+1.5\,u(k-9)u(k)+0.1,$
where $y(k)$ is the output sequence and $u(k)$ is a scalar driving input (often i.i.d. uniform in [0, 0.5]00.5[0,\,0.5][ 0 , 0.5 ]). In this one-step-ahead prediction task, the reservoir receives $u(k)$ at time $k$ and must produce $\hat{y}(k+1)$ that approximates the true next value $y(k+1)$. The main challenge is that $y(k+1)$ depends on its previous ten outputs in a highly nonlinear manner, requiring the reservoir to capture temporal dependencies spanning at least ten steps.
IV-A3 Discussion
Fig. 3 summarizes the MC reservoir’s predictive performance on these tasks under different simulation conditions (rows) and time horizons (columns). The leftmost column illustrates the reservoir’s one-step-ahead forecasts for the NARMA10 series; the remaining three columns correspond to Mackey–Glass predictions at horizons of 1, 10, and 20 steps, respectively. Within each panel, the true time series (blue) is overlaid with the predicted values (red).
Rows (a–d) in Fig. 3 show the deterministic mean-field simulation results. Rows (e–h) show particle-based stochastic simulations that capture molecular noise. We observe that although the more realistic stochastic simulations introduce additional fluctuations in the reservoir node states, the overall predictive accuracy remains quite comparable to the deterministic case, indicating that the approximate mean-field analysis is valid and that the fluctuations do not severely degrade performance.
Focusing on the Mackey–Glass predictions (columns 2–4), we see that the reservoir accurately tracks the chaotic trajectory for modest horizons (e.g., $\Delta=1$ or $\Delta=10$), but it can still capture essential dynamics even at $\Delta=20$. Unsurprisingly, the error grows at larger horizons, yet the predictions remain qualitatively coherent with the true signal. This capability underscores the MC reservoir’s capacity for handling long-range memory in a strongly nonlinear system.
In the leftmost column, the reservoir’s one-step predictions for NARMA10 closely match the target sequence. Although NARMA10 typically requires the memory of up to ten past outputs, the reservoir’s nonlinear fading dynamics prove sufficient to encode that history and generate accurate forecasts.
IV-B Optimization of the Reservoir
IV-B1 Effect on MC System Parameters on Performance
The computational efficacy of the proposed molecular reservoir computer is intricately related to the biophysical parameters that govern the MC channel. Specifically, the dynamics of ligand-receptor binding, molecular propagation, and symbol encoding collectively define the temporal complexity and nonlinearity accessible to the reservoir. To systematically evaluate these dependencies, we conducted a parametric study involving some key variables: the ligand-receptor binding and unbinding rate parameters governing the receiver response {$k_{on}$, $k_{off}$}, the symbol interval and input amplitude governing transmitted signal{$N_{max}$, $T$}, and the propagation channel parameters such as diffusion coefficient (D) and transmitter-receiver distance $x_{0}$. The effects of changing these parameters in a physically plausible range are presented in Fig. 4 through the metric of NRMSE.
Our findings reveal a nonmonotonic relationship between parameter settings and predictive performance, as measured by the NRMSE. Specifically, the interplay between binding dynamics (i.e., the ratio $k_{off}$/$k_{on}$), inter-symbol duration, diffusion rate, and transmitter-receiver distance plays a critical role in determining the channel’s computational effectiveness. Low $k_{off}$/$k_{on}$ ratios, short inter-symbol durations, and high diffusion coefficients—especially when combined with shorter distances—lead to receiver saturation. In these scenarios, the occupation ratio of the receptor remains consistently high, nearing full saturation with minimal temporal variation. As a result, the system exhibits reduced dynamical richness and a diminished capacity to encode time-varying input features, which leads to an increase in NRMSE and degraded predictive accuracy. This effect is particularly evident in high-density ligand environments where excessive receptor occupation eliminates the temporal diversity necessary for reservoir computation.
Conversely, conditions characterized by high $k_{off}$/$k_{on}$ ratios or insufficient diffusion coefficients impede the ability of ligands to effectively bind and form receptor-ligand complexes. In such settings, the MC channel fails to facilitate reliable signal transmission, leading to sparse or inconsistent occupation patterns. This not only reduces the input-driven modulation of the reservoir state but also undermines the system’s ability to perform the nonlinear transformation essential for projecting input signals into a higher-dimensional computational space. Moreover, the consistent trends in prediction performance observed across both the NARMA10 and Mackey–Glass benchmarks reinforce the generalizability of these parameter–performance relationships, demonstrating that the interplay between binding kinetics and diffusion dynamics fundamentally governs the temporal richness and nonlinearity of the reservoir, and thus provides clear guidelines for optimizing MC‑based reservoir computers under diverse biochemical conditions.
Figure 5: Effect of the moving average filter window length on the NRMSE for stochastic simulations.
IV-B2 Post-Process Filtering
Due to the inherent stochasticity of molecular diffusion and receptor-ligand interactions, especially in low-concentration regimes, the raw output of the molecular reservoir exhibits noise-driven fluctuations, as observed in Fig. 2. These fluctuations, though biologically realistic, can obfuscate the underlying information-carrying trends. To address this, we applied a moving average filter to the received signal of receptor occupation $b(t)$, effectively attenuating high-frequency noise components while preserving meaningful temporal dynamics. The entire received signal sequence is filtered, the effects of which on the prediction task are shown in Fig. 3. The post-processed received signals are observed to display outputs similar to those of the estimation layer to deterministic simulations, enabling more robust performance during inference. The change in error with respect to the length of the filtering window is presented in Fig. 5, the filter improves performance until a certain point, after which the loss of information becomes too great and leads to an increase in error.
V Conclusion
In this work, we have demonstrated, through cross‑validated deterministic and particle‑based stochastic simulations, that a canonical point‑to‑point MC channel naturally realizes a physical reservoir computer. By interpreting the time‑multiplexed receptor occupancy as virtual nodes, the system reliably forecasts Mackey–Glass and NARMA10 time series with low normalized error, despite substantial molecular noise. A comprehensive parameter sweep analysis revealed that computational power of the MC reservoir hinges on a delicate balance between several system parameters, including receptor binding kinetics, diffusion coefficients, and symbol intervals, all of which influence the memory and nonlinearity essential for effective reservoir computing. This analysis provides the initial guidelines for rationally optimizing MC systems for wetware AI. Future work will focus on extending the system to MC networks and multiple ligand-receptor pairs, and on implementing the proposed physical reservoir computer in experimental MC platforms.

### v2

@@
 Abstract
-Molecular Communication (MC) channels inherently possess significant memory and nonlinear dynamics due to diffusion and receptor kinetics, often posing challenges for reliable data transmission.
-This work reconceptualizes these intrinsic properties as computational resources by framing a canonical point-to-point MC channel, consisting of ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a physical reservoir computer (PRC).
-We utilize the time-varying fraction of bound receptors as the reservoir’s internal state, employing time-multiplexing to generate high-dimensional virtual nodes without explicit recurrence.
-Only a linear readout layer is trained via ridge regression.
-Through deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s capability for complex temporal processing by successfully performing next-step prediction on standard chaotic time-series benchmarks (Mackey-Glass and NARMA10).
-Performance, quantified by Normalized Root Mean Square Error (NRMSE), exhibits a non-monotonic dependence on key system parameters (receptor kinetic rates, diffusion coefficient, transmitter-receiver distance), revealing optimal operational regimes.
-These findings validate the potential of using MC channel as effective and low-complexity computational substrate.
-Index Terms:
-Reservoir Computing, Molecular Communications, Ligand-Receptor Interactions
+Molecular Communication (MC) channels are characterized by significant memory and nonlinear dynamics arising from diffusion and receptor kinetics.
+While often viewed as impairments to reliable data transmission, this work introduces a paradigm shift by reconceptualizing these intrinsic physical properties as computational resources.
+We frame a canonical point-to-point MC channel, comprising ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a Physical Reservoir Computer (PRC).
+Utilizing deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s inherent capability for complex temporal information processing on standard chaotic time-series benchmarks.
+We comprehensively evaluate performance using both task-specific Normalized Root Mean Square Error (NRMSE) and the task-independent Information Processing Capacity (IPC).
+Our results reveal a non-monotonic dependence of computational power on key biophysical parameters (receptor kinetic rates, diffusion coefficient, and transmitter-receiver distance), identifying optimal operational regimes where memory and nonlinearity are balanced.
+These findings establish the MC channel as a viable computational substrate, paving the way for novel architectures in wetware artificial intelligence.
 I Introduction
 Reservoir Computing (RC), introduced independently as the echo‑state network (ESN) by Jaeger [1] and the liquid‑state machine (LSM) by Maass et al.
-[2], is a class of recurrent neural network models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression.
-This separation of representation and learning obviates back‑propagation through time, providing fast, data‑efficient training while retaining the rich temporal processing power of recurrent dynamics.
-As a result, RC has proven highly effective for time-series prediction tasks [3].
-Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implements large virtual networks through time-multiplexing of a single nonlinear node [8, 9].
+[2], is a class of recurrent neural network (RNN) models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression.
+The separation of representation and learning in RC bypasses the need for back-propagation through time, enabling fast, data‑efficient training while retaining the computational power of recurrent dynamics.
+As a result, RC has proven highly effective for challenging time-series prediction tasks [3].
+Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implement large virtual networks through time-multiplexing of a single nonlinear node [8, 9].
 Such implementations greatly reduce hardware complexity while achieving performance on par with digital RC, and have catalyzed interest in neuromorphic edge computing [10] and embodied intelligence [11].
-Recent work has begun to evaluate biological systems as reservoirs, capitalizing on their inherently rich and adaptive dynamics.
+Recent work has evaluated biological systems as reservoirs, capitalizing on their inherently rich, adaptive, and energy-efficient dynamics.
 Pioneering experiments with neuron cultures [12], brain organoids [13], bacteria [14], and plants [15] demonstrate that living systems can be co-opted or integrated into artificial computing frameworks to execute externally defined computational tasks.
@@
 Building on these naturally evolved mechanisms, researchers have long studied engineered molecular communication (MC) systems as alternative to conventional electromagnetic (EM) communications to enable networks of micro/nanoscale devices via the exchange of precisely controlled chemical signals [18].
-This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19].
+This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19, 20].
 Extensive theoretical research and recent experimental work in this field have particularly focused on overcoming the unique challenges of MC channels that are different than EM communications, including significant channel memory effects (inter-symbol interference) and nonlinear receiver responses (involving receptor saturation) in addition to biochemical noise [21, 22].
@@
 In this paper, we reconceptualize MC channel as a physical reservoir computer, revealing its potential for information processing and implementing machine learning tasks.
-Specifically, we study a simple MC channel model, which is adequately representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors.
-By adopting a time-multiplexed RC framework, we leverage the inherent channel memory arising from diffusion and the nonlinear ligand-receptor interactions at the receiver, both of which are critical for computational performance.
-Through numerical analysis and particle-based spatial stochastic simulations, we demonstrate that MC channels can successfully perform chaotic time-series prediction tasks with minimal training, as benchmarked by standard RC evaluations on NARMA10 and Mackey-Glass time-series.
-We systematically investigate how various MC system parameters, including the number of transmitted molecules, transmitter-receiver distance, diffusion coefficient, and ligand-receptor kinetic rates, affect computing performance.
-Our results provide evidence that even simple MC channels can be harnessed for non-trivial time-series tasks by virtue of their intrinsic diffusion-driven memory and receptor-mediated nonlinearities.
+Specifically, we study a canonical MC channel model, which is representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors.
+We demonstrate that the inherent physical dynamics of this communication system, namely the channel memory from ligand diffusion and the nonlinearity from receptor kinetics, can be directly leveraged for complex computation.
+This is achieved by adopting a time-multiplexed RC framework to translate these continuous physical states into a high-dimensional computational state space with minimal overhead.
+To characterize the computational power of the MC reservoir, we conduct a comprehensive parametric study, quantifying performance through two complementary lenses.
+First, we investigate task-specific predictive accuracy on standard chaotic benchmarks (Mackey–Glass and NARMA10).
+Second, we evaluate its task-independent computational power by analyzing the resulting Information Processing Capacity (IPC).
+This dual analysis provides critical insights as to how performance is fundamentally governed by key biophysical parameters, such as ligand-receptor kinetic rates, diffusion coefficient, and transmission distance, generating the first guidelines for optimizing MC systems for computation.
+Our results demonstrate that even simple MC channels possess the intrinsic capacity for non-trivial time-series processing.
 Beyond contributing to the fundamental understanding of information-processing capabilities of biological systems at the elemental level of their communication links, this computing perspective on MC channels can open new avenues for bio-inspired and bio-integrated computing architectures.
-Potential applications range from early disease diagnosis analyzing time-varying biomarker profiles to channel equalization in artificial MC networks and IoBNT [22].
-By demonstrating how the channel’s inherent dynamics can be systematically exploited for information processing, this study lays the groundwork for MC systems to serve not only as physical channels for communication but also as powerful computational substrates in their own right.
+Potential applications range from real-time analysis of time-varying biomarker profiles for early disease diagnosis to channel equalization in artificial MC networks and IoBNT [23].
+By demonstrating how the channel’s inherent dynamics can be exploited for information processing, this study lays the groundwork for MC channels to serve not only as physical medium for unconventional communications but also as powerful computational substrates in their own right.
 The paper is organized as follows.
-Section II reviews molecular communication and physical reservoir computing.
-Section III formalizes the molecular‑communication channel as a physical reservoir computer and describes the modeling and training methodology.
-Section IV presents deterministic and stochastic simulation results with parameter studies.
-Section V concludes and outlines future directions.
+Section II reviews the fundamentals of MC and PRC.
+Section III formalizes the MC channel as a physical reservoir computer and describes the modeling and training methodology.
+Section IV presents the performance evaluation, using both task-specific NRMSE and task-independent IPC to reveal the optimal operational regimes governed by the MC system parameters.
+Section V concludes the study and outlines future directions.
 II Background
 II-A Molecular Communications
-Bio-inspired MC emerged as a promising approach to realize networks of micro/nanoscale devices, such as biosensors and microrobots, and to seamlessly interface these engineered networks with biological systems.
-Such integration underpins the vision of the IoBNT, where MC mediates collaborative tasks such as targeted drug delivery and continuous health monitoring [23].
-By exploiting mechanisms similar to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods become impractical [24].
-From early theoretical studies on capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [21, 25, 26, 27].
+Bio-inspired MC has emerged as a key enabling technology for realizing networks of micro/nanoscale devices, such as biosensors and microrobots, and seamlessly interfacing them with biological systems [18].
+This integration underpins the IoBNT vision, where MC mediates collaborative tasks such as smart drug delivery and continuous health monitoring [24].
+By exploiting mechanisms analogous to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods are impractical [25].
+From early theoretical studies on channel capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [22, 26, 27, 28].
 Researchers have identified unique challenges that distinguish MC from EM communications, including severe inter-symbol interference (ISI) due to slow diffusion, and nonlinearity plus memory effects from stochastic receptor binding and other reaction kinetics.
 Practical MC transceivers at the microscale also face tight constraints in energy, size, and computational resources, motivating new strategies in encoding, detection, and synchronization [21].
-To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [28].
+To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [29, 30].
 In parallel, an expanding line of research investigates MC for in situ artificial intelligence (AI) with living cells and engineered biomolecular processes, commonly referred to as wetware AI.
@@
 Another study integrated biologically plausible spiking neurons into reservoir computing, using MC-based connectivity for distributed processing among cells [33].
-A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GNRs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [32].
+A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GRNs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [34].
 By contrast, bottom-up efforts explicitly engineer MC as the foundation for novel neural network architectures.
@@
 II-B Physical Reservoir Computing
-In a standard RC scheme, the input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
+In a standard RC scheme (Fig. 1(a)), an input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
 $\mathbf{x}(t+1)=f\!\bigl(\mathbf{A}\,\mathbf{x}(t)+\mathbf{W}_{\mathrm{in}}\mathbf{u}(t)\bigr),$ (1)
-where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinearity (Fig.
-1(a)) [36].
-The readout is obtained by a linear map
-$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t),$ (2)
-with $\mathbf{W}_{\mathrm{out}}$ trained, often by ridge regression, to minimize the discrepancy between $\mathbf{y}(t)$ and a target signal.
-Because the internal weights $\mathbf{A}$ remain fixed, RC sidesteps the heavy cost of back‑propagating gradients through deep recurrent structures, a feature that makes it well‑suited to hardware realizations.
+where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ is the input weight matrix that projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinear activation function [38].
+The readout is generated by a linear transformation of the reservoir state:
+$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t).$ (2)
+The readout weights $\mathbf{W}_{\mathrm{out}}$ are trained, typically via ridge regression, to minimize the error between $\mathbf{y}(t)$ and a target signal.
+Since $\mathbf{A}$ and $\mathbf{W}_{\mathrm{in}}$ remain fixed, RC sidesteps the computationally intensive gradient descent required for training traditional RNNs, making it highly suitable for hardware implementation.
 PRC replaces the software‑defined reservoir with the intrinsic dynamics of a physical substrate.
@@
 1(b)) [9, 37].
-Leveraging the medium’s native parallelism, bandwidth, and energy efficiency, PRC offers a neuromorphic alternative to digital simulations.
-Time‑multiplexed reservoirs.
-High‑dimensional states can be generated with minimal hardware by time‑multiplexing a single nonlinear element (Fig.
+By exploiting the medium’s native parallelism and energy efficiency, PRC offers a powerful neuromorphic alternative to digital computation.
+A key innovation in PRC is the use of time-multiplexing to generate high-dimensional reservoir states with minimal hardware, often utilizing just a single nonlinear node (Fig.
 1(c)) [8].
-During each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants.
+In this approach, during each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants.
 The resulting vector
 $\mathbf{x}(t)=\bigl[x_{1}(t),x_{2}(t),\dots,x_{M}(t)\bigr]^{\!\top}$
-acts as an $M$‑node virtual reservoir whose evolution from one period to the next is governed by the masked input and the device’s nonlinear response [37].
-A widely adopted form of time‑multiplexing employs an external feedback loop of delay $\tau$ around a single nonlinear node (e.g., a semiconductor laser or an analogue circuit).
-Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ yields $M$ virtual nodes [9, 38].
-Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal supplies fading memory that underpins tasks such as speech recognition and chaotic time‑series forecasting.
+constitutes an MM‑node virtual reservoir. The evolution of the reservoir state is governed by the interaction of the masked input and the reservoir’s inherent nonlinear dynamics and memory [39].
+A common implementation of time-multiplexing involves an external feedback loop with delay $\tau$ around the nonlinear node (e.g., a semiconductor laser or an analogue circuit).
+Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ defines the $M$ virtual nodes [9, 40].
+Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal provides the necessary fading memory for tasks such as speech recognition and chaotic time‑series forecasting.
 Additional delay lines or delayed taps at the input or readout can further enrich the dynamics or extend the memory horizon without increasing the number of physical components [41].
@@
 1(d).
-The transmitter release a prescribed number of ligands (molecules) instantaneously at prearranged times into a fluidic medium, where they propagate by free diffusion.
-A fraction of these ligands eventually arrives near the receiver and can bind reversibly to receptors on its surface under given reaction kinetics.
-Below, we describe how this widely studied MC system model can be reframed as a physical reservoir computer by regarding the time-varying fraction of bound receptors as the high-dimensional dynamical state.
-Let the transmitter release $I(n)$ molecules at each symbol time $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval.
-For simplicity, we assume the receiver is at distance $x_{R}\gg r_{\mathrm{rx}}$ from the transmitter at the origin, such that the receiver’s finite radius $r_{\mathrm{rx}}$ can be neglected in the channel impulse response.
-Under three-dimensional (3D) free diffusion with diffusion coefficient $D$, the ligand concentration $c(\mathbf{x},t)$ obeys Fick’s second law:
+The transmitter releases a specific number of ligands (molecules) instantaneously at predefined intervals into a fluidic medium.
+These ligands propagate via free diffusion, and a fraction eventually reaches the receiver, where they bind reversibly to surface receptors under given reaction kinetics.
+We now detail how this canonical MC system is formalized as a physical reservoir computer, utilizing the time-varying fraction of bound receptors as the dynamical state.
+Let the transmitter release $I(n)$ molecules at discrete symbol times $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval.
+We assume a 3D environment where the ligand concentration $c(\mathbf{x},t)$ evolves according to Fick’s second law of diffusion:
 $\frac{\partial c(\mathbf{x},t)}{\partial t}\;=\;D\,\nabla^{2}c(\mathbf{x},t),$ (3)
-with the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$. The impulse response at distance $x_{0}$ for a single pulse released at $t=0$ is the 3D diffusion kernel
-$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl{(}-\frac{x_{0}^{2}}{4Dt}\Bigr{)}\quad(t>0).$ (4)
-Hence, when the transmitter emits $I(n)$ molecules at times $nT$, the local ligand concentration at the receiver (treated as a point in the far-field approximation) is given by
-$c_{T}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl{(}t-nT\bigr{)}\,u\bigl{(}t-nT\bigr{)},$ (5)
-where $u(\cdot)$ is the unit step function enforcing causality.
-Since $h(t)$ has a infinitely long tail (though it decays rapidly), successive transmissions introduce ISI.
-For practical numerical simulations, this tail is often truncated to a finite channel memory length [40].
+where $D$ is the diffusion coefficient of ligands. We impose the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$.
+Assuming the receiver is fixed at a distance $d$ significantly larger than its radius $r_{\mathrm{rx}}$ ($d\gg r_{\mathrm{rx}}$), we can approximate the receiver as a point when calculating the channel impulse response.
+The impulse response at distance $d$ for a pulse released at $t=0$ is then given by the 3D diffusion kernel:
+$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl(-\frac{d^{2}}{4Dt}\Bigr)\quad(t>0).$ (4)
+The local ligand concentration at the receiver, $c_{\mathrm{R}}(t)$, resulting from the sequence of transmissions $I(n)$ is the superposition:
+$c_{\mathrm{R}}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl(t-nT\bigr).$ (5)
+The long tail of the diffusion kernel $h(t)$ inherently introduces ISI, providing the channel with signigicant memory.
+In numerical simulations, this tail is often truncated to a finite memory length [42].
 Figure 2: (a) An instance from particle-based stochastic simulations of the MC channel with Smoldyn.
-Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey-Glass series input, (c) NARMA10 Series input, (d) Mackey-Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
-The spherical receiver is equipped with $N_{R}$ identical, monovalent ligand receptors. Each receptor can be either unbound or bound to one ligand, with reversible reaction:
+Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey–Glass series input, (c) NARMA10 series input, (d) Mackey–Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
+The spherical receiver is equipped with $N_{R}$ identical, monovalent receptors.
+The interaction between ligands ($L$) and unbound receptors ($R$) is modeled as a reversible reaction forming a ligand-receptor complex ($LR$):
 $L+R\overset{k_{\text{on}}}{\underset{k_{\text{off}}}{\longleftrightarrow}}LR,$
-where $L$ is the free ligand, $R$ is an unbound receptor, and $LR$ is a ligand-receptor complex, i.e., bound receptor.
-At the mean-field level, and assuming the local free ligand concentration is not significantly depleted by bound receptors, the time evolution of the mean fraction of bound receptors, $b(t)=n_{b}(t)/N_{R}$, can be approximated by
-$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{T}(t)\,\bigl{[}1-b(t)\bigr{]}\;-\;k_{\mathrm{off}}\,b(t),$ (6)
-where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the forward (binding) and reverse (unbinding) rate constants, respectively, and $c_{T}(t)$ is the local ligand concentration at the receiver.
-Although the exact time-varying variance of $b(t)$ under such nonstationary input is typically not available in closed form (due to correlated ligand arrivals and receptor occupancy fluctuations), we use particle-based spatial stochastic simulations (with Smoldyn (Fig.
-2) [41]) in Section IV to account for this stochasticity.
-For the present analytical treatment, (6) provides a deterministic approximation to the mean behavior of the ligand-receptor interactions.
-[…]
-In RC, one aims to transform an input time series $\{u(n)\}$ into a high-dimensional dynamical state $\mathbf{x}_{n}$ through a fixed nonlinear system (the reservoir), and then train only a linear readout for the target computational task.
-Here, we map the input $u(n)$ to the number of released molecules $I(n)$ at each symbol time $nT$:
+At the mean-field level, assuming the local ligand concentration is not significantly depleted by binding, the time evolution of the mean fraction of bound receptors, $b(t)=n_{\mathrm{b}}(t)/N_{\mathrm{R}}$, can be approximated by
+$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{\mathrm{R}}(t)\,\bigl[1-b(t)\bigr]\;-\;k_{\mathrm{off}}\,b(t),$ (6)
+where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the association (binding) and dissociation (unbinding) rate constants, respectively.
+This equation captures the essential nonlinearity (due to the saturation term [1b(t)1-b(t)]) and the memory (due to the kinetics) of the receiver response.
+Equation (6) provides a deterministic, mean-field approximation of the receptor dynamics.
+This model is computationally efficient and allows for extensive parametric analysis.
+However, MC systems are inherently stochastic due to the discrete nature of molecules and the probabilistic nature of diffusion and binding events.
+To validate the mean-field model and assess the impact of molecular noise, we also employ particle-based spatial stochastic simulations using the Smoldyn simulator (Fig.
+2) [43] in Section IV.
+Fig.
+2(a) shows a snapshot from a Smoldyn simulation, illustrating the diffusion of ligands from the point source to the spherical receiver.
+Figs.
+2(b) and 2(c) compare the evolution of the receptor occupation ratio $b(t)$, i.e., the reservoir’s internal state, for the deterministic model and the stochastic simulation when driven by two standard chaotic time series (Mackey–Glass and NARMA10).
+The stochastic simulations exhibit significant noise.
+However, when a moving-average filter ($2000$ ms window) is applied (Figs.
+2(d) and 2(e)), it is evident that the deterministic model accurately captures the mean behavior of the stochastic system, justifying its use for analyzing the fundamental computational dynamics.
+In the RC framework, the input time series $\{u(n)\}$ must be mapped into a high-dimensional, nonlinear dynamical state $\mathbf{x}_{n}$.
+We encode the input into the number of released molecules at time $nT$:
 $I(n)\;=\;\alpha\,u(n),$ (7)
-where $\alpha$ is a scaling factor. The channel then imposes a complex, memory-rich temporal response on these inputs (due to both diffusion and receptor binding).
-To extract a sufficiently high-dimensional state from the scalar output $b(t)$, we perform time-multiplexing.
-That is, within each interval $[nT,\,(n+1)T]$, we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$:
+where $\alpha$ is a scaling factor. The MC channel naturally imposes a complex, memory-rich transformation on this input sequence, resulting in the continuous-time scalar output $b(t)$.
+To construct a high-dimensional reservoir state from $b(t)$, we employ time-multiplexing.
+Within each symbol interval [nT,(n+1)T][nT,\,(n+1)T], we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$, yielding the reservoir state vector:
 $\mathbf{x}_{n}\;=\;\bigl[b\bigl(nT+\tau_{1}\bigr),\;b\bigl(nT+\tau_{2}\bigr),\;\dots,\;b\bigl(nT+\tau_{M}\bigr)\bigr]^{\top}.$ (8)
-This sampling effectively creates $M$ virtual nodes per symbol.
-Unlike a typical optoelectronic delay-based reservoir, however, we do not require a separate external feedback loop; rather, the channel’s intrinsic ISI provides the dynamical memory.
-The “masking” function typically used in conventional time-multiplexed RC is replaced here by the MC channel impulse response, which shapes the input in time to eventually generate diverse node responses upon sampling.
-We train a linear readout to map from the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a time-series prediction).
-Let $\mathbf{W}_{\mathrm{out}}$ be the $M$-dimensional readout weight vector, then
+This process creates $M$ virtual nodes per symbol, akin to time-multiplexing used in delay-based RC, where a single nonlinear node is sampled at multiple, equally spaced intervals to emulate a high-dimensional network.
+In a typical delay-based RC, an external mask is required to diversify the dynamics of the virtual nodes, and an explicit delay loop provides memory.
+In the MC reservoir, these requirements are fulfilled intrinsically by the physical channel.
+The diffusion process inherently reshapes each molecular pulse in time, acting as a natural, dynamic mask.
+The resulting ISI from diffusion, combined with the memory from receptor kinetics, provides the necessary fading memory.
+Moreover, the receptor (receiver response) nonlinearity ensures that the temporal samples are dynamically rich and diverse.
+The final step involves training a linear readout to map the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a future value of the time series).
+The prediction $\hat{y}_{n}$ is given by:
 $\hat{y}_{n}\;=\;\mathbf{W}_{\mathrm{out}}^{\!\top}\,\mathbf{x}_{n}.$ (9)
-With $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, we solve:
+Given $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, the readout weights $\mathbf{W}_{\mathrm{out}}$ are optimized using ridge regression to minimize the regularized mean square error:
 $\min_{\mathbf{W}_{\mathrm{out}}}\Bigl\|\mathbf{y}\;-\;\mathbf{X}\,\mathbf{W}_{\mathrm{out}}\Bigr\|^{2}\;+\;\lambda\,\|\mathbf{W}_{\mathrm{out}}\|^{2},$ (10)
-where $\mathbf{X}$ is the $N_{\mathrm{train}}\!\times\!M$ matrix whose $n$-th row is $\mathbf{x}_{n}^{\top}$, $\mathbf{y}$ is the vector of target outputs $y_{n}$, and $\lambda$ is a regularization parameter (ridge regression).
-The closed-form solution is then given by the Moore–Penrose pseudoinverse:
+where $\mathbf{X}$ is the state matrix (whose $n$-th row is $\mathbf{x}_{n}^{\top}$), $\mathbf{y}$ is the target vector, and $\lambda$ is a regularization parameter.
+The closed-form solution is given by the Moore–Penrose pseudoinverse:
 $\mathbf{W}_{\mathrm{out}}\;=\;\bigl(\mathbf{X}^{\top}\,\mathbf{X}\;+\;\lambda\,\mathbf{I}\bigr)^{\!-1}\,\mathbf{X}^{\top}\,\mathbf{y}.$ (11)
-Hence, only $\mathbf{W}_{\mathrm{out}}$ is optimized, and the underlying channel and binding dynamics remain unmodified.
+Importantly, only $\mathbf{W}_{\mathrm{out}}$ is trained, while the underlying physical dynamics of the MC channel remain fixed.
 IV Results
-In this section, we present the results of the performance evaluation for the MC reservoir computer using two complementary approaches.
-First, deterministic analyses were carried out in MATLAB using the mean‑field kinetic model (Eq.
-(6)) with the default simulation settings listed in Table I.
-Second, to capture the stochasticity of the MC channel, we performed particle‑based spatial stochastic simulations in Smoldyn [41].
-By comparing the outcomes of the deterministic and stochastic simulations, we both confirm the validity of the approximate deterministic model and quantify the effects of intrinsic noise on RC performance.
+Here, we present a comprehensive performance evaluation for the MC reservoir computer.
+The analysis is structured to first demonstrate the MC reservoir’s core capabilities, then to investigate the factors that govern its computational power.
+We begin by benchmarking the performance on two standard chaotic time-series prediction tasks: the Mackey–Glass and NARMA10 series.
+We investigate the impact of inherent molecular noise on the performance and demonstrate how post-process low-pass filtering can be used to improve prediction accuracy.
+We then conduct a comprehensive parametric analysis to reveal how key MC system parameters, such as receptor kinetics, diffusion rate, and transmitter-receiver distance, can be tuned to optimize performance.
+This analysis is carried out through two complementary metrics: the task-specific NRMSE and the task-independent IPC.
+Unless otherwise specified, all analyses utilize the default parameters listed in Table I, wherein the parameters defining the physical MC system are chosen to represent physiologically relevant values consistent with established literature in MC [21, 22].
 TABLE I: Default Values of Simulation and RC Parameters
 Parameter Value
-Receptor binding constant $k_{\text{on}}$ $6.022\times 10^{8}\,\mathrm{M^{-1}s^{-1}}$
-Receptor unbinding constant $k_{\text{off}}$ $1\,\mathrm{s^{-1}}$
 Maximum # released molecules $N_{\max}$ $3000$
@@
 Symbol duration $T$ $1\,\mathrm{s}$
-Mackey–Glass prediction length $\Delta$ $6$ steps
 Diffusion coefficient $D$ $1\times 10^{-11}\,\mathrm{m^{2}s^{-1}}$
-Tx–Rx distance $x_{0}$ $10\,\upmu\mathrm{m}$
-Reservoir nodes $M$ $100$
-Numerical (deterministic) soln. timestep $\Delta t_{d}$ $1\,\mathrm{ms}$
-Stochastic simulation timestep $\Delta t_{s}$ $10\,\mathrm{ms}$
-Stochastic simulation receiver radius $R$ $3\,\upmu\mathrm{m}$
+Transmitter-receiver distance $d$ $10\,\upmu\mathrm{m}$
+Receiver radius $R$ $3\,\upmu\mathrm{m}$
+Receptor binding constant $k_{\text{on}}$ $6.022\times 10^{8}\,\mathrm{M^{-1}s^{-1}}$
+Receptor unbinding constant $k_{\text{off}}$ $1\,\mathrm{s^{-1}}$
+Reservoir virtual nodes $M$ $100$
+Mackey–Glass prediction length $\Delta$ 66 steps
 Training sequence length $2000$
 Test sequence length $2000$
-IV-A Chaotic Time-Series Prediction
-We benchmark our MC reservoir computer on two widely used chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [42].
-These tasks have become standards for assessing both algorithmic and physical RC architectures because they combine complex temporal dependencies, strong nonlinearities, and a requirement for long‑range memory.
-The goal is to learn a mapping from the present (and past) states of the time series to future values, using only a linear readout layer trained on the reservoir states.
-Performance Metric: We quantify the predictive accuracy of the reservoir using the normalized root-mean-square error (NRMSE),
-$\text{NRMSE}\;=\;\sqrt{\frac{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\hat{y}_{k}\,]^{2}}{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\bar{y}\,]^{2}}},$
-where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples.
-An NRMSE closer to 00 indicates a more accurate prediction.
-Figure 4: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey-Glass series, and (d-f) for NARMA10 series prediction tasks.
-IV-A1 Mackey-Glass Series
-The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
+Numerical simulation timestep $\Delta t_{\mathrm{d}}$ $1\,\mathrm{ms}$
+Stochastic simulation timestep $\Delta t_{\mathrm{s}}$ $10\,\mathrm{ms}$
+IV-A Performance on Benchmark Time-Series Prediction Tasks
+We benchmark the MC reservoir computer on two widely adopted chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [44].
+These benchmarks are standards for evaluating RC architectures as they demand a combination of complex temporal dependency modeling, strong nonlinearity, and significant memory depth.
+IV-A1 Computing Tasks and Performance Metric
+Mackey–Glass Series: The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
 $\frac{dx(t)}{dt}=\beta\,\frac{x(t-\tau)}{1+x(t-\tau)^{n}}-\gamma\,x(t),$ (12)
@@
 For the canonical parameter set $\beta=0.2$, $\gamma=0.1$, $n=10$, and $\tau=17$, the system exhibits chaotic behavior with a fractal attractor.
-We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation using a suitable step size (e.g., $0.1$), then sampling at discrete time points $k$.
+We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation, then sampling at discrete time points $k$.
 The reservoir input at time step $k$ is set to $u(k)=x_{k}$, and the task is to predict $x_{k+\Delta}$, i.e., the value of the series $\Delta$ steps into the future.
-IV-A2 NARMA10 Series
-A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by
+NARMA10 Series: A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by the recursive equation:
 $y(k+1)$ $=0.3\,y(k)\;+\;$ (13)
@@
 where $y(k)$ is the output sequence and $u(k)$ is a scalar driving input (often i.i.d.
-uniform in [0, 0.5]00.5[0,\,0.5][ 0 , 0.5 ]).
+uniform in [0, 0.5][0,\,0.5]).
 In this one-step-ahead prediction task, the reservoir receives $u(k)$ at time $k$ and must produce $\hat{y}(k+1)$ that approximates the true next value $y(k+1)$.
 The main challenge is that $y(k+1)$ depends on its previous ten outputs in a highly nonlinear manner, requiring the reservoir to capture temporal dependencies spanning at least ten steps.
-IV-A3 Discussion
+Figure 3: Prediction performance of the MC reservoir on two chaotic time-series benchmarks.
+Columns: (left) one–step prediction of the NARMA10 sequence; (center–left, center–right, right) prediction of the Mackey–Glass sequence at prediction horizons of 1, 10, and 20 symbols, respectively.
+Rows: (a–d) deterministic mean‑field analysis; (e–h) particle‑based stochastic simulation; (i–l) stochastic simulation post‑processed with a $2000\,\text{ms}$ moving‑average filter.
+In every panel the blue curve is the ground‑truth target and the red curve is the reservoir prediction.
+Performance Metric: Predictive accuracy is quantified using NRMSE:
+$\text{NRMSE}\;=\;\sqrt{\frac{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\hat{y}_{k}\,]^{2}}{\sum_{k=1}^{N_{\mathrm{test}}}[\,y_{k}-\bar{y}\,]^{2}}},$
+where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples.
+Lower NRMSE indicates higher accuracy.
+IV-A2 Time-Series Prediction Results
 Fig.
-3 summarizes the MC reservoir’s predictive performance on these tasks under different simulation conditions (rows) and time horizons (columns).
-The leftmost column illustrates the reservoir’s one-step-ahead forecasts for the NARMA10 series; the remaining three columns correspond to Mackey–Glass predictions at horizons of 1, 10, and 20 steps, respectively.
-Within each panel, the true time series (blue) is overlaid with the predicted values (red).
-Rows (a–d) in Fig.
-3 show the deterministic mean-field simulation results.
-Rows (e–h) show particle-based stochastic simulations that capture molecular noise.
-We observe that although the more realistic stochastic simulations introduce additional fluctuations in the reservoir node states, the overall predictive accuracy remains quite comparable to the deterministic case, indicating that the approximate mean-field analysis is valid and that the fluctuations do not severely degrade performance.
-Focusing on the Mackey–Glass predictions (columns 2–4), we see that the reservoir accurately tracks the chaotic trajectory for modest horizons (e.g., $\Delta=1$ or $\Delta=10$), but it can still capture essential dynamics even at $\Delta=20$.
-Unsurprisingly, the error grows at larger horizons, yet the predictions remain qualitatively coherent with the true signal.
-This capability underscores the MC reservoir’s capacity for handling long-range memory in a strongly nonlinear system.
-In the leftmost column, the reservoir’s one-step predictions for NARMA10 closely match the target sequence.
-Although NARMA10 typically requires the memory of up to ten past outputs, the reservoir’s nonlinear fading dynamics prove sufficient to encode that history and generate accurate forecasts.
-IV-B Optimization of the Reservoir
-IV-B1 Effect on MC System Parameters on Performance
-The computational efficacy of the proposed molecular reservoir computer is intricately related to the biophysical parameters that govern the MC channel.
-Specifically, the dynamics of ligand-receptor binding, molecular propagation, and symbol encoding collectively define the temporal complexity and nonlinearity accessible to the reservoir.
-To systematically evaluate these dependencies, we conducted a parametric study involving some key variables: the ligand-receptor binding and unbinding rate parameters governing the receiver response {$k_{on}$, $k_{off}$}, the symbol interval and input amplitude governing transmitted signal{$N_{max}$, $T$}, and the propagation channel parameters such as diffusion coefficient (D) and transmitter-receiver distance $x_{0}$.
-The effects of changing these parameters in a physically plausible range are presented in Fig.
-4 through the metric of NRMSE.
-Our findings reveal a nonmonotonic relationship between parameter settings and predictive performance, as measured by the NRMSE.
-Specifically, the interplay between binding dynamics (i.e., the ratio $k_{off}$/$k_{on}$), inter-symbol duration, diffusion rate, and transmitter-receiver distance plays a critical role in determining the channel’s computational effectiveness.
-Low $k_{off}$/$k_{on}$ ratios, short inter-symbol durations, and high diffusion coefficients—especially when combined with shorter distances—lead to receiver saturation.
-In these scenarios, the occupation ratio of the receptor remains consistently high, nearing full saturation with minimal temporal variation.
-As a result, the system exhibits reduced dynamical richness and a diminished capacity to encode time-varying input features, which leads to an increase in NRMSE and degraded predictive accuracy.
-This effect is particularly evident in high-density ligand environments where excessive receptor occupation eliminates the temporal diversity necessary for reservoir computation.
-Conversely, conditions characterized by high $k_{off}$/$k_{on}$ ratios or insufficient diffusion coefficients impede the ability of ligands to effectively bind and form receptor-ligand complexes.
-In such settings, the MC channel fails to facilitate reliable signal transmission, leading to sparse or inconsistent occupation patterns.
-This not only reduces the input-driven modulation of the reservoir state but also undermines the system’s ability to perform the nonlinear transformation essential for projecting input signals into a higher-dimensional computational space.
-Moreover, the consistent trends in prediction performance observed across both the NARMA10 and Mackey–Glass benchmarks reinforce the generalizability of these parameter–performance relationships, demonstrating that the interplay between binding kinetics and diffusion dynamics fundamentally governs the temporal richness and nonlinearity of the reservoir, and thus provides clear guidelines for optimizing MC‑based reservoir computers under diverse biochemical conditions.
+3 summarizes the MC reservoir’s predictive performance on the NARMA10 and Mackey–Glass tasks across three simulation settings: a deterministic mean-field model (a-d), a particle-based stochastic model capturing molecular noise (e-h), and the stochastic model with a post-processing filter applied to the reservoir state (i-l).
+The deterministic simulations (Fig.
+3a-d) provide a performance baseline, revealing the computational potential of the idealized MC reservoir.
+The reservoir achieves high-level accuracy on the one-step-ahead Mackey–Glass prediction (MG-1, NRMSE = 0.054), demonstrating its ability to capture short-term temporal dependencies.
+For longer horizons, we observe a non-trivial trend: performance degrades for MG-10 (NRMSE = 0.354) but then counter-intuitively improves for MG-20 (NRMSE = 0.311).
+We attribute this improvement not to a simple monotonic decay of memory, but to the specific autocorrelation function (ACF) of the Mackey–Glass series for $\tau=17$.
+The ACF for this chaotic series is quasi-periodic, exhibiting decaying oscillations (see, e.g., Fig.
+2(b) in [41]).
+Our results suggest that for the given sampling rate, the prediction horizon of $\Delta=10$ steps corresponds to a lag time where the ACF magnitude is near a local minimum, i.e., close to a zero-crossing.
+At such a point, the linear correlation between the present state and the future target is weakest, making the prediction task inherently challenging for a linear readout layer and resulting in a higher NRMSE.
+In contrast, the $\Delta=20$ horizon appears to align with a region of resurgent correlation, i.e., a secondary peak in the ACF’s magnitude.
+Although this correlation is weaker than at very short lags, its greater magnitude compared to the lag at $\Delta=10$ provides the linear readout with a more stable and informative target, leading to a lower NRMSE.
+Therefore, the performance of the MC-PRC on chaotic time-series prediction is modulated not just by the reservoir’s memory decay, but also by the intrinsic correlation structure of the time-series itself.
+The introduction of molecular noise in the particle-based stochastic simulations (Fig.
+3e-h) expectedly degrades performance, yet reveals more complex dynamics such that the one-step-ahead Mackey–Glass prediction (MG-1) becomes significantly less accurate than the 20-step-ahead prediction (NRMSE of 0.865 vs.
+0.702, respectively).
+We propose this counter-intuitive result stems from the relationship between the amplitude of the signal to be predicted and the inherent noise floor of the physical reservoir.
+For the MG-1 task, the target represents a subtle, incremental change in the smoothly-sampled time series.
+The stochastic fluctuations in the reservoir state, driven by discrete molecular events, are of a comparable magnitude and can therefore obscure this fine-grained signal from the linear readout.
+In contrast, the MG-20 prediction task targets a much larger-amplitude change due to the system’s chaotic divergence over a longer horizon.
+This larger signal remains clearly discernible above the reservoir’s intrinsic noise floor, allowing the readout to learn a more robust mapping and achieve a lower error.
+To mitigate the detrimental effects of this noise, we apply a $2000$ ms moving-average filter to the reservoir’s raw output state before the readout stage (Fig.
+3i-l).
+This post-processing step yields substantial improvements in predictive accuracy across all stochastic scenarios, with NRMSE values approaching those of the deterministic case (e.g., for MG-10, improving from $0.921$ to $0.583$).
+The quantitative effect of the filter window length, detailed in Fig.
+4, reveals that the optimal filtering strategy is highly task-dependent.
+For the Mackey–Glass series, the NRMSE decreases monotonically with increasing window length before plateauing, suggesting its useful signal is encoded in low-frequency dynamics that benefit from aggressive noise removal.
+In stark contrast, the NARMA10 series exhibits a distinct U-shaped performance curve with an optimal window around $1000$ ms, indicating that essential, higher-frequency information would be erased by an overly long filter.
+Interestingly, this filtering alters the performance trend for Mackey–Glass predictions visible in Fig.
+3: the MG-20 task (NRMSE = $0.653$) is now slightly less accurate than the MG-10 task (NRMSE = $0.583$).
+This reversal suggests that while the filter effectively removes detrimental high-frequency noise, its smoothing effect may also attenuate the subtle, low-frequency periodicities in the reservoir state that the unfiltered system exploited for the improved 20-step-ahead prediction.
 Figure 4: Effect of the moving average filter window length on the NRMSE for stochastic simulations.
-IV-B2 Post-Process Filtering
-Due to the inherent stochasticity of molecular diffusion and receptor-ligand interactions, especially in low-concentration regimes, the raw output of the molecular reservoir exhibits noise-driven fluctuations, as observed in Fig.
-2.
-These fluctuations, though biologically realistic, can obfuscate the underlying information-carrying trends.
-To address this, we applied a moving average filter to the received signal of receptor occupation $b(t)$, effectively attenuating high-frequency noise components while preserving meaningful temporal dynamics.
-The entire received signal sequence is filtered, the effects of which on the prediction task are shown in Fig.
-3.
-The post-processed received signals are observed to display outputs similar to those of the estimation layer to deterministic simulations, enabling more robust performance during inference.
-The change in error with respect to the length of the filtering window is presented in Fig.
-5, the filter improves performance until a certain point, after which the loss of information becomes too great and leads to an increase in error.
+IV-B Parametric Analysis for Reservoir Optimization
+We now conduct a comprehensive parametric analysis to understand the physical principles governing the MC reservoir’s performance and to identify its optimal operating regimes.
+We evaluate how key biophysical parameters influence the reservoir’s computational power using our two complementary metrics: the task-specific NRMSE and the task-independent IPC.
+To render this extensive parameter exploration computationally tractable, all subsequent analyses in this section are performed using the deterministic mean-field model.
+Figure 5: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey–Glass series, and (d-f) for NARMA10 series prediction tasks.
+IV-B1 Effect of MC System Parameters on Task Performance
+The computational efficacy of the proposed MC reservoir is intricately governed by the biophysical parameters of the underlying communication channel.
+To elucidate these relationships, we systematically evaluated how performance on our benchmark tasks, quantified by NRMSE, depends on key physical parameters.
+The results, presented in Fig.
+5, reveal clear operational regimes and trade-offs that are fundamental to designing such systems.
+Receiver Kinetics
+Figs.
+5(a) and (d) explore the interplay between the receptor binding ($k_{\mathrm{on}}$) and unbinding ($k_{\mathrm{off}}$) constants.
+For both the Mackey–Glass and NARMA10 tasks, optimal performance (lowest NRMSE) is confined to a distinct diagonal region.
+This reveals that performance is not dictated by either rate alone, but by their ratio, which determines the receptor’s binding affinity (inversely related to the dissociation constant, $K_{\mathrm{D}}=k_{\mathrm{off}}/k_{\mathrm{on}}$).
+Two failure regimes are evident.
+At very low $k_{\mathrm{on}}$ or high $k_{\mathrm{off}}$ (low affinity), the system enters a sparse-binding regime where too few molecules are captured to form a reliable signal.
+Conversely, while a very low $k_{\mathrm{off}}$ (high affinity) indeed increases the reservoir’s memory by extending receptor occupancy time, this comes at the cost of dynamical responsiveness.
+In this regime, the reservoir state becomes dominated by its own long-timescale memory, effectively locking in past information.
+The influence of new, incoming molecular signals is consequently suppressed, as they are unable to significantly modulate the already-occupied receptor population.
+This loss of sensitivity to the input signal diminishes the reservoir’s ability to perform the nonlinear transformations necessary for computation.
+Therefore, optimal performance requires a finely tuned intermediate binding affinity.
+This balance provides sufficient memory to capture relevant temporal dependencies while preserving the crucial ability for the reservoir state to be dynamically and richly modulated by the present input.
+Signal Encoding and Transmission Rate
+Figs.
+5(b) and (e) reveal that the optimal configuration of the number of released molecules ($N_{\mathrm{max}}$) and symbol duration ($T$) is fundamentally task-dependent.
+In our system, both parameters co-determine the reservoir’s two essential computational ingredients: fading memory and nonlinearity.
+The symbol duration $T$ directly tunes the structure of the memory by controlling ISI, while both $T$ and $N_{\mathrm{max}}$ jointly determine the effective molecular concentration at the receiver, which dictates the operating point on its nonlinear response curve.
+For the Mackey–Glass task (Fig.
+5(b)), which requires predicting the evolution of a smooth, autonomous trajectory, the optimal region exhibits a diagonal trend where increasing $T$ requires a compensatory increase in $N_{\mathrm{max}}$.
+This suggests a need to maintain a consistent level of nonlinear activation; as pulses become more temporally isolated, they must be stronger to induce a comparable dynamical response.
+In contrast, the NARMA10 task (Fig.
+5(e)) shows a markedly different behavior where performance is more strongly dependent on $N_{\mathrm{max}}$ than on $T$.
+Across all symbol durations beyond the initial high-ISI regime ($T>1$ s), performance consistently improves as $N_{\mathrm{max}}$ is decreased.
+We attribute this to the specific computational demands of the NARMA10 task, which, unlike the trajectory-following MG task, requires the nonlinear processing of a sequence of distinct, random input values.
+Channel Propagation
+Figs.
+5(c) and (f) map performance across the diffusion coefficient ($D$) and transmitter-receiver distance ($d$), which together sculpt the channel’s impulse response and govern the signal strength that drives the receiver’s nonlinearity.
+The results reveal a striking diagonal band of optimal performance, demonstrating that computational success hinges not on independently tuning these parameters, but on achieving a critical balance between the memory timescale and the strength of nonlinear activation.
+In the regime of low $D$ and high $d$, severe signal attenuation weakens the arriving molecular pulse, causing the receiver to operate in a near-linear, computationally poor regime despite its long memory.
+Conversely, at high $D$ and low $d$, the characteristic transport time, scaling with $d^{2}/D$, becomes vanishingly short.
+The channel approaches a memoryless limit, and although the receiver’s nonlinearity is strongly engaged, it lacks the temporal context necessary for prediction.
+The optimal diagonal thus represents the sweet spot where the channel’s memory kernel is appropriately structured for the task, while the signal remains potent enough to drive the system into a computationally rich, nonlinear regime.
+IV-B2 Information Processing Capacity (IPC) Analysis
+To evaluate the computational capability of the MC reservoir more fundamentally, we employ the IPC framework introduced by Dambre et al.
+[45], and further refined by Schulte-to-Brinke et al.
+[46].
+IPC quantifies how effectively the system states can approximate various nonlinear and memory-dependent transformations of past inputs, providing a comprehensive, task-independent measure of computational richness.
+Formally, IPC is computed by projecting the input history onto an orthonormal polynomial basis—specifically, the set of Legendre polynomials $\{P_{n}(\cdot)\}$, as they provide a complete orthonormal basis for signals defined on the interval [1,1][-1,1].
+Given an input signal $u(t)$, normalized to have zero mean and constrained within the interval [1,1][-1,1], we define polynomial functionals of the form:
+$y(t)=\prod_{i}P_{n_{i}}\left(u(t-s_{i})\right),$ (14)
+where $n_{i}$ is the polynomial order and $s_{i}$ is the delay associated with the $i^{\text{th}}$ factor.
+Each unique set $\{(n_{i},s_{i})\}$ is termed a factor set and represents a specific computational task defined by polynomial order and memory depth.
+We systematically enumerate factor sets using an ordered backtracking algorithm to ensure uniqueness and completeness while avoiding redundant evaluations [46].
+For each factor set, we compute the optimal linear estimator $\hat{y}(t)$ from the system states $\mathbf{X}(t)$ via ridge regression (using pseudo-inverse methods) to avoid numerical instability.
+The corresponding capacity $C$ for each factor set is then defined as the squared correlation coefficient between the actual target $y(t)$ and its linear estimator $\hat{y}(t)$:
+$C=\text{Corr}^{2}\left(y(t),\hat{y}(t)\right).$ (15)
+Figure 6: Comparison of IPC contributions for different order and delay of univariate Legendre terms.
+Figure 7: Parametric analysis of the MC reservoir’s IPC.
+The color in each panel represents the total IPC, summed over all significant first- and second-order Legendre polynomial terms.
+(a) IPC as a function of the receiver kinetics, governed by the binding ($k_{\mathrm{on}}$) and unbinding ($k_{\mathrm{off}}$) constants.
+(b) IPC as a function of the signal encoding, determined by the maximum number of released molecules ($N_{\mathrm{max}}$) and the symbol duration ($T$).
+(c) IPC as a function of the channel propagation properties, defined by the diffusion coefficient ($D$) and the transmitter-receiver distance.
+Due to finite-sample estimation effects, very small IPC values may occur purely by chance.
+To mitigate these spurious capacities, we employ a statistical significance threshold based on the $\chi^{2}$-distribution as detailed in [45, 46].
+Specifically, we select a significance threshold corresponding to a false-positive rate $p=10^{-4}$, and all IPC values below this threshold are set to zero.
+Total IPC is then the sum of all significant individual capacities:
+$\text{IPC}_{\text{total}}=\sum_{\text{significant }C}C,$ (16)
+thereby providing a rigorous measure of the system’s overall computational power.
+Finally, IPC results are presented either as total capacities, or subdivided by polynomial order, delay, or specific factor sets, facilitating a detailed interpretation of system performance.
+Our analysis of IPC uncovers meaningful trends in how both polynomial order and memory delay influence capacity within the univariate Legendre polynomial basis (Fig.
+6).
+As expected for any RC system, increasing the memory delay leads to a systematic reduction in IPC contribution.
+This reflects the inherent challenge of precisely reconstructing information from the distant past in a physical system, where signals naturally decay and become increasingly corrupted by temporal dispersion.
+Likewise, higher-order polynomials consistently yield lower IPC contributions.
+This is because more complex nonlinear transformations demand finer, more precise mappings from the reservoir state, which are inherently more difficult for a physical system with limited computational power to realize.
+Given the prominent roles played by first and second-order polynomial terms in relevant computational tasks, such as NARMA10 and Mackey–Glass time series prediction, our subsequent parametric analysis in Fig.
+7 specifically focuses on the combined IPC of these orders.
+Fig.
+7(a) illustrates the effect of ligand–receptor interaction rates ($k_{\mathrm{on}}$ and $k_{\mathrm{off}}$) on IPC.
+The analysis reveals a well-defined optimal region at intermediate rates, surrounded by three distinct low-IPC regimes that correspond to the failure to achieve the necessary balance between memory and nonlinearity.
+We identify the following computationally poor regimes.
+First, at very low $k_{\mathrm{on}}$ and very high $k_{\mathrm{off}}$, the system is both memoryless and weakly coupled to the input; ligands bind rarely and unbind almost instantly, preventing the reservoir from either retaining past information or responding strongly to the present input.
+Second, in the regime of very low $k_{\mathrm{on}}$ and very low $k_{\mathrm{off}}$, the reservoir is effectively inert; the weak input coupling is combined with a static memory, meaning the state, once changed, fails to evolve in response to new information.
+The third and most critical failure mode occurs at very high $k_{\mathrm{on}}$ and very low $k_{\mathrm{off}}$.
+Here, the MC system enters deep receptor saturation.
+The high binding rate and long occupancy time cause the reservoir state to lock in, losing its dynamic range and sensitivity to input fluctuations.
+While this configuration provides a long memory, it is a static, computationally poor memory that suppresses the rich, graded nonlinear transformations essential for high IPC.
+The optimal region is therefore the sweet spot that successfully balances these competing factors.
+Fig.
+7(b) plots the IPC across the number of released molecules ($N_{\mathrm{max}}$) and the symbol duration ($T$).
+Consistent with our task-specific NRMSE analysis, the IPC is severely degraded at very short T due to the formation of unstructured memory, where catastrophic ISI destroys the temporal resolution necessary for computation.
+Beyond this initial high-ISI regime, IPC becomes largely independent of $T$ and is instead governed mainly by $N_{\mathrm{max}}$.
+This behavior is reminiscent of our findings for the NARMA10 task, suggesting that IPC (constructed with first and second-order terms) is more informative of the prediction performance in NARMA10 task compared to MG task.
+Fig.
+7(c) illustrates the IPC across the diffusion coefficient ($D$) and transmitter-receiver distance ($d$).
+A quasi-diagonal band of optimal IPC is observed, highlighting that computational richness is achieved by tuning the channel’s physical properties to match a characteristic timescale.
+In the region of low $D$ and high $d$, IPC is low.
+Here, signal attenuation is severe, leading to a weak signal at the receiver.
+This fails to activate the receiver’s nonlinearity, causing the system to operate in a near-linear, computationally poor regime.
+Although the memory kernel is long, its quality is poor due to high temporal dispersion and low signal strength.
+These findings reveal the necessity of jointly tuning both physical and temporal system parameters to fully exploit the computational capacity of MC-based reservoir.
+The trends observed in the IPC results demonstrate a meaningful relationship with the reservoir’s empirical performance on time series prediction.
+Although IPC and NRMSE are distinct metrics, i.e., one measuring fundamental capacity, the other task-specific performance, their close behaviors across parameter sweeps make IPC a valuable theoretical tool for anticipating favorable operating conditions and guiding the design and tuning of MC-based reservoir computers.
+However, the precise quantitative link between a high total IPC and high performance on a specific task remains an open and challenging research question in the reservoir computing field [47].
 V Conclusion